OpenStudy (popstar30977):

JOGGING Jaren jogs at a rate of 7.5 miles per hour. How many miles per minute does jarin jog? (please put the example of the problem)

OpenStudy (popstar30977):

@satellite73

OpenStudy (anonymous):

how many minutes in an houjr?

OpenStudy (popstar30977):

60

OpenStudy (anonymous):

divide

OpenStudy (mertsj):

\[\frac{7.5 mi}{1 hr}\times\frac{1hr}{60\min}=\frac{7.5 mi}{60 \min}=.125 \frac{mi}{\min}\]

OpenStudy (popstar30977):

is thatt all i have to do or do i have to multiply 1hr x 60min x 60 min @Mertsj

OpenStudy (mertsj):

Look at the problem setup. The units cancel just like numbers so the hr on top and the hr on the bottom cancel.

OpenStudy (popstar30977):

ok thanks alot @Mertsj

OpenStudy (mertsj):

Then on top you have 7.5mi and on bottom you have 60 min.

OpenStudy (mertsj):

Divide7.5by 60 and get .125 with the units mi/min

OpenStudy (mertsj):

yw

OpenStudy (popstar30977):

the answer is 28.8 right @Mertsj

OpenStudy (mertsj):

Do you have a calculator?